# Their currents turn awry, and lose the name of action. I: Fundamental limits to orbit reconstruction due to non-conservation of stellar actions

### v1

@@ 4.2 Change in actions over time
and similarly for $z$, where here $D$ is the diffusion coefficient; we carry out this fit from $\Delta t=5-50$ Myr for $D_{R}$ and for $5-40$ Myr for $D_{z}$, covering the age range when the behaviour is close to linear, and for the purpose of the fit we include all stars regardless of age $t_{*}$. We show our fits, together with the data, in Figure 9. Our best-fitting values for the diffusion coefficients are $D_{R}=2.5$ Gyr-1 and $D_{z}=4.8$ Gyr-1. The inverses of these values correspond to characteristic diffusion timescales of approximately 400 Myr for $J_{R}$ and 200 Myr for $J_{z}$, meaning that over these timescales individual stars forget their individual actions. The shorter timescale for $J_{z}$ reflects stronger dynamical perturbations in the vertical direction. In both cases, however, we see that the curve in Figure 8 flattens beyond $\approx 100$ Myr, indicating that actions have reached maximum decorrelation by this point, corresponding to around 50% median square relative change. By this age stellar actions are essentially drawn at random from the distribution of actions for stars of the appropriate age, and retain no memory of their values $\approx 100$ Myr earlier.
Figure 9: Same as Figure 8, but now zooming in on ages $<100$ Myr. The red dashed lines indicate linear fits to the data, as described in the main text. The expressions and $R^{2}$ values for the fits are indicated in the legend, with times in units of Myr.
@@ 4.5 Comparison with initial stars
As described in subsection 2.2, the WK23 simulations include a population of stellar disc particles that are already present in the initial conditions of Z25. We refer to these pre-existing particles as ‘initial stars’. In contrast, the star particles discussed so far are those born during the simulation and hence are all younger than 464 Myr. To understand the underlying mechanisms driving action evolution, it is illuminating to compare these two populations. Since the potential grid used to calculate action (see section 3 and Appendix A) extends only to the disc, this comparison is restricted to stars within the disc region.
The key differences between these two populations are their vertical height distributions and their distribution relative to dense gas. Initial stars have positions that are uncorrelated with gas, while stars that form in the simulation are necessarily in close proximity to dense gas structures at least at birth. Similarly, the root mean square height $\langle z^{2}\rangle^{1/2}$ of initial stars remains between $380-405$ pc during the simulation, whereas, as already mentioned, for stars born during the simulation, it is much lower, between $130-200$ pc. Thus the initial stars spend much more time away from the midplane while their younger counterparts remain confined to the midplane.
In Figure 12, we show the median square relative changes in radial and vertical actions, $\delta J_{R}^{2}$ and $\delta J_{Z}^{2}$, as a function of time lag for stars of different stellar ages born in the simulation (same as in Figure 8) along with the same for initial stars, represented by the blue dashed line. Despite the differences between the two populations, the change in radial action over time is quite similar. The fact that even the oldest stars experience a comparable level of radial action change to newly born stars suggests that this change is driven by a mechanism that is present throughout the disc and does not depend strongly on birth conditions and proximity to small-scale gas structures. The natural candidate that meets this requirements are transient spiral arms, which are known to be a key feature of disc dynamics and which extend over a range of vertical heights, since for a Milky Way-like disc they involve strongly-coupled perturbations in both gas and stars (e.g., Romeo & Mogotsi, 2017).
@@
To further confirm our hypotheses, we examine phase-space orbits of sample initial stars and stars born in the simulation in Figure 13. We select for this purpose stars whose action evolution lies close to the median values of their respective populations, ensuring that these orbits are typical trajectories for both the categories. Since we applied a low-pass filter to the action time series (see subsection 3.3), we handpick stars that do not seem to be part of clusters to minimise the impact of filtering. The figure shows that orbits do confirm the expected trend: initial stars with larger actions have nearly-closed orbits, as expected for a conserved action, while the stars born during the simulation follow more irregular trajectories, which is consistent with stronger vertical action evolution for the younger population.
@@ 5 Conclusion
  • The distributions of actions over our simulation time of $\approx 500$ Myr remains almost constant, broadening only slightly. For parts of the disc interior to $\approx 12$ kpc, and thus interior to the warp in the Milky Way disc, we find good agreement between the rate of broadening in the simulation and that observed in the Milky Way.
@@
While the rapid action evolution we measure might at first seem surprising, it is important to remember that many previous studies of stellar action conservation have been conducted in the context of a purely collisionless stellar disc, and even in these cases actions are not perfectly conserved (e.g., Solway et al., 2012; Mikkola et al., 2020). Other studies that include gas find that its gravitational influence strongly scatters young stars (e.g., Fujimoto et al., 2023), and the general importance of gas becomes apparent if we recall that, while gas makes up only $\sim 15\%$ of the mass of the Milky Way disc, its smaller scale height means that at the midplane gas and stars contribute about equally to the vertical gravitational acceleration (e.g., McKee et al., 2015; Krumholz et al., 2018). In this context it is worth considering as an example the Local Bubble, which is effectively a void in the local ISM, which otherwise has a mean density of $\sim 1$ atom cm-3, several hundred pc across (Linsky et al., 2022; Zucker et al., 2022; O’Neill et al., 2024). The free-fall time of this structure is (Equation 2) $\approx 51.5$ Myr, comparable to the epicyclic periods in the Solar neighbourhood ($T_{\kappa}\approx 170$ Myr and $T_{\nu}\approx 85$ Myr; Binney & Tremaine, 2008), and its size means stars passing through this region experience a non-axisymmetric gravitational perturbations over a significant portion of their orbit. It is therefore not surprising that a structure of this size should be able to perturb stellar orbits significantly.
@@
Our results imply that disc stars, unlike halo stars, do not follow regular, predictable orbits, and therefore that reliable stellar orbit reconstruction relative to the disc past $\sim 100$ Myr is not feasible. However, this does not necessarily mean that star clusters are entirely irrecoverable. While individual stars undergo a random walk in logarithmic action space, our preliminary analysis indicates that stars born together tend to remain correlated with each other in action space over time. This correlation is not surprising, given that much of the change in action appears to be driven by relatively large-scale perturbations such as the Local Bubble – stars born closer to one another than the characteristic sizes of such structures are subjected to similar perturbations, and therefore should experience comparable changes in their actions and remain together in action space. In essence, while stars drift randomly in action space, those that form together “hold hands” as they move, thus preserving clustering signatures in action space that could still be used to reconstruct star clusters and associations, if not to trace them back to their formation locations. We will explore this idea further in Paper II of this series.
@@ Acknowledgements
This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI Australia) and the Pawsey Supercomputing Centre, NCRIS-enabled capabilities supported by the Australian Government, through award jh2. AA and MRK acknowledge support from the Australian Research Council through Laureate Fellowship FL220100020. MI wishes to thank HW Rix for lunchtime discussions in 2019 that motivated some of this work, and the Australia-Germany Joint Research Co-operation Scheme of Universities Australia for funding this travel.
@@
The codes and a subset of selected intermediate data used in this study are available at https://github.com/aruuniima/stellar-actions-I and https://www.mso.anu.edu.au/~arunima/stellar-actions-I-data/. The initial simulation outputs and potential fits are not included in the public repository due to their size, but will be provided by the authors upon reasonable request.
Appendix A Computing epicyclic frequencies
@@
In this appendix, we describe in detail the procedure that we use for calculating epicyclic frequencies $\kappa$ and $\nu$ for each star particle. We begin with the 2D potential grid that spans from 0.1 pc to 20 kpc in the radial direction and from $-1$ to $+1$ kpc in the vertical direction, computed as described in subsection 3.1, and we must evaluate partial derivatives of this potential to compute $\kappa$ and $\nu$. Naive numerical evaluation of these derivatives yields extremely noisy estimates, and we must therefore adopt a smoothing procedure.
@@
The guiding radius $R_{g}$ and radial epicyclic frequency depend on $\partial\Phi/\partial R$ and $\partial^{2}\Phi/\partial R^{2}$ evaluated at $z=0$. Our first step to estimate these quantities is to extract the potential grid values within ±$\pm 2$ pc of the midplane, corresponding to five grid points in the vertical direction, and average them to produce a one-dimensional function of $R$. This averaging reduces the noise. We next apply a Gaussian filter [^1] This and all subsequent operations use the implementations provided in SciPy version 1.11.4; see Virtanen et al. (2020). with a standard deviation of 30 pc using ‘nearest’ mode to prevent artificial edge effects; this further smooths the data. Finally, we fit the smoothed data with a B-spline with smoothing parameter $s=200$ to obtain a continuous, differentiable function $\Phi(R,z=0)$.
@@
With this smooth function at hand, we are now prepared to evaluate the radial derivatives of $\Phi$ and the quantities that depend upon them. We first define a new radial grid with a resolution of 0.5 pc with its inner edge placed at the larger of 2 kpc and 5 kpc inward from the smallest radial position of any star at the time we are sampling, and its outer edge at the smaller of 17.5 kpc and 5 kpc outward from the largest stellar radial position; this ensures that we have good coverage over the full disc, but that we are not extrapolating to radii where we sample the potential poorly. We then evaluate $\Omega_{i}$ at each radial grid point $R_{i}$ from Equation 5 using our B-spline representation of $\Phi(R)$. The corresponding specific angular momentum at each grid point is $L_{z,i}=R_{i}^{2}\Omega_{i}$, and thus our grid represents a set of $(R_{i},L_{z,i})$ pairs. Since the guiding radius $R_{g}$ for each star is defined implicitly by its angular momentum $L_{z,*}$ and the condition $L_{z,*}=R_{g}^{2}\Omega(R_{g})$, we can use the $(R_{i},L_{z,i})$ pairs as a lookup table to find the guiding radius corresponding to any $R_{g}$. In practice we do this by defining a new B-spline function $R_{g}(L_{z})$ from our $(R_{i},L_{z,i})$ pairs using a smoothing factor $s=2$, and use the resulting function to evaluate $R_{g}$ for every star. We then drop from our sample stars for which $R_{g}$ lies outside the range $2-16$ kpc, on the grounds that they lie too close to our grid edge to be reliable.
@@
The calculation of the vertical epicyclic frequency $\nu$ follows a similar but slightly more complex process, since we require the second derivative of $\Phi(z)$ at each star’s guiding radius (see Equation 7), rather than for the radial case which only requires derivatives of $\Phi(R)$ evaluated at $z=0$. To achieve this, we construct a sparse radial grid with a resolution of 200 pc, spanning the range $2-16$ kpc to match the guiding radius constraints that we imposed earlier. At each radial grid point, we extract the potential from the 2D grid by averaging over all radial values within ±$\pm 1$ pc of the sparse grid point. This averaging, as in the radial case, reduces local noise and provides a better estimate of the galactic potential as a function of $z$. We then smooth the resulting $\Phi(z)$ data by applying a Savitzky-Golay filter with a window length of 101 pc and a polynomial order of 2, which smooths the data while preserving the underlying curvature, followed by a Gaussian filter with standard deviation of 2 pc, which we apply three times. Finally, we use the smoothed data as input to a B-spline fit with $s=50$ to generate a smooth, continuously-differentiable functional form of $\Phi(z)$ at each sparse radial grid point. We use this smooth function to evaluate $\partial^{2}\Phi/\partial z^{2}$ at $z=0$, the quantity required to evaluate $\nu$ using Equation 7. This yields a set of $(R_{i},\nu_{i})$ pairs on our sparse radial grid. We fit these data with a B-spline fit with $s=1000$ to obtain a continuous function $\nu(R)$. Finally, we determine $\nu$ for each star by inputting its guiding radius into this function.

### v2

@@ 5.1 Change in actions over time
and similarly for $z$, where here $D$ is the diffusion coefficient; we carry out this fit from $\Delta t=5-50$ Myr for $D_{R}$ and for 5405-40 Myr for $D_{z}$, covering the age range when the behaviour is close to linear, and for the purpose of the fit we include all stars regardless of age $t_{*}$. We show our fits, together with the data, in Figure 11. Our best-fitting values for the diffusion coefficients are $D_{R}=2.5$ Gyr-1 and $D_{z}=4.8$ Gyr-1. The inverses of these values correspond to characteristic diffusion timescales of approximately 400 Myr for $J_{R}$ and 200 Myr for $J_{z}$, meaning that over these timescales individual stars forget their individual actions. The shorter timescale for $J_{z}$ reflects stronger dynamical perturbations in the vertical direction. In both cases, however, we see that the curve in Figure 10 flattens beyond $\approx 100$ Myr, indicating that actions have reached maximum decorrelation by this point, corresponding to around 50% median square relative change. By this time, stellar actions are essentially drawn at random from the distribution of actions for stars of the appropriate age, and retain no memory of their values $\approx 100$ Myr earlier.
Note that the timescale that we measure here for changes in the actions of individual stars is much shorter (at least an order of magnitude) than the timescale for evolution of the population-median action or population velocity dispersion we find in Section 4. This difference arises because the two approaches measure fundamentally different processes. Population-level statistics such as AVRs or action distributions are sensitive only to changes across all stars, so for example if the actions of two stars were swapped these statistics would remain unchanged. By contrast, our absolute and relative action-change metrics defined in equation 16 and equation 17, even when summarised via their median over a large number of stars, would register a large change in such a swap. This explains how the population median action can require Gyr to evolve even while individual stars change their actions on a much smaller timescale as we see here. It also highlights that stability in the distribution of actions does not necessarily imply conservation of actions for individual stars.
Figure 11: Same as Figure 10, but now zooming in on time lag $\Delta t$ <100<100 Myr. The red dashed lines indicate linear fits to the data, as described in the main text. The expressions and $R^{2}$ values for the fits are indicated in the legend, with times in units of Myr.
@@ 5.4 Comparison with initial stars
As described in Section 2.2, the WK23 simulations include a population of stellar disc particles that are already present in the initial conditions of Z25. We refer to these pre-existing particles as ‘initial stars’. In contrast, the star particles discussed so far are those born during the simulation and hence are all younger than 464 Myr. To understand the underlying mechanisms driving action evolution, it is illuminating to compare these two populations. Since the potential grid used to calculate action (see Section 3 and Appendix B) extends only to the disc, this comparison is restricted to stars within the disc region.
The key differences between these two populations are their vertical height distributions and their distribution relative to dense gas. Initial stars have positions that are uncorrelated with gas, while stars that form in the simulation are necessarily in close proximity to dense gas structures at least at birth. Similarly, the root mean square height $\langle z^{2}\rangle^{1/2}$ of initial stars remains between 380405380-405 pc during the simulation, whereas, as already mentioned, for stars born during the simulation, it is much lower, between 130200130-200 pc. Thus the initial stars spend much more time away from the midplane while their younger counterparts remain confined to the midplane.
In Figure 14, we show the median square relative changes in radial and vertical actions, $\delta J_{R}^{2}$ and $\delta J_{Z}^{2}$, as a function of time lag for stars of different initial stellar ages born in the simulation (same as in Figure 10) along with the same for initial stars, represented by the blue dashed line. Despite the differences between the two populations, the change in radial action over time is quite similar. The fact that even the oldest stars experience a comparable level of radial action change to newly born stars suggests that this change is driven by a mechanism that is present throughout the disc and does not depend strongly on birth conditions and proximity to small-scale gas structures. The natural candidate that meets this requirements are transient spiral arms, which are known to be a key feature of disc dynamics and which extend over a range of vertical heights, since for a Milky Way-like disc they involve strongly-coupled perturbations in both gas and stars (e.g., Romeo & Mogotsi, 2017).
@@
To further confirm our hypotheses, we examine phase-space orbits of sample initial stars and stars born in the simulation in Figure 15. We select for this purpose stars whose action evolution lies close to the median values of their respective populations, ensuring that these orbits are typical trajectories for both the categories. Since we applied a low-pass filter to the action time series (see LABEL:subsec:_action_evolution), we handpick stars that do not seem to be part of clusters to minimise the impact of filtering. The figure shows that orbits do confirm the expected trend: initial stars with larger actions have nearly-closed orbits, as expected for a conserved action, while the stars born during the simulation follow more irregular trajectories, which is consistent with stronger vertical action evolution for the younger population.
@@ 6 Conclusion
  • The distributions of actions over our simulation time of $\approx 500$ Myr remains almost constant, broadening only slightly. For parts of the disc interior to $\approx 12$ kpc, and thus interior to the warp in the Milky Way disc, we find good agreement between the rate of broadening in the simulation and that observed in the Milky Way. We also find that the age-velocity relation in our simulations is relatively flat, with only modest increases in the stellar velocity dispersion over the simulation time.
@@
While the rapid action evolution we measure might at first seem surprising, it is important to remember that many previous studies of stellar action conservation have been conducted in the context of a purely collisionless stellar disc, and even in these cases actions are not perfectly conserved (e.g., Solway et al., 2012; Mikkola et al., 2020). Other studies that include gas find that its gravitational influence strongly scatters young stars (e.g., Fujimoto et al., 2023), and the general importance of gas becomes apparent if we recall that, while gas makes up only $\sim 15\%$ of the mass of the Milky Way disc, its smaller scale height means that at the midplane gas and stars contribute about equally to the vertical gravitational acceleration (e.g., McKee et al., 2015; Krumholz et al., 2018). In this context it is worth considering as an example the Local Bubble, which is effectively a void in the local ISM (interstellar medium), which otherwise has a mean density of $\sim 1$ atom cm-3, several hundred pc across (Linsky et al., 2022; Zucker et al., 2022; O’Neill et al., 2024). The free-fall time of this structure is (equation 2) $\approx 51.5$ Myr, comparable to the epicyclic periods in the Solar neighbourhood ($T_{\kappa}\approx 170$ Myr and $T_{\nu}\approx 85$ Myr; Binney & Tremaine, 2008), and its size means stars passing through this region experience a non-axisymmetric gravitational perturbations over a significant portion of their orbit. It is therefore not surprising that a structure of this size should be able to perturb stellar orbits significantly.
Our results imply that disc stars, unlike halo stars, do not follow regular, predictable orbits, and are subject to more rapid and erratic action evolution. Even in the halo, orbit reconstruction is difficult: Arora et al. (2022) find that even in the absence of major mergers, integration of orbits in a time-independent potential renders orbit reconstruction of stellar streams unreliable after only 0.5–1 Gyr. This inability of orbit recovery is only likely to be more severe in the disc, where dynamical times are shorter and perturbations from gas and substructure are more frequent and intense. This idea is further supported by Kamdar et al. (2021), who study stellar streams in a dynamical model of Milky Way-like disc and find that even with perfect information about the non-axisymmetric structures in the galaxy (spiral arms and a bar), the fraction of stream stars that could be reliably integrated backward in time drops below half within 100 Myr. With integration in only axisymmetric potential, the fraction drops below half within few tens of Myr. Notably, their work uses an analytical potential with idealised structure, whereas our simulation includes live ISM, self-consistent star formation and MHD physics. Therefore, in the more realistic galactic environment of our simulation, we conclude that reliable stellar orbit reconstruction relative to the disc past $\sim 100$ Myr is not feasible.
@@
However, this does not necessarily mean that star clusters are entirely irrecoverable. While individual stars undergo a random walk in logarithmic action space, our preliminary analysis indicates that stars born together tend to remain correlated with each other in action space over time. This correlation is not surprising, given that much of the change in action appears to be driven by relatively large-scale perturbations such as the Local Bubble – stars born closer to one another than the characteristic sizes of such structures are subjected to similar perturbations, and therefore should experience comparable changes in their actions and remain together in action space. In fact, the errors introduced by non-conservation of actions are likely to be similar for stars that form close together, meaning that while their individual orbits may be inaccurate, backward integration will still place them near the same location – enabling recovery of the cluster, though not its orbit relative to the Galactic disc. In essence, while stars drift randomly in action space, those that form together “hold hands” as they move, thus preserving clustering signatures in action space that could still be used to reconstruct star clusters and associations, if not to trace them back to their formation locations. We will explore this idea further in Paper II of this series.
@@ Acknowledgements
We thank the anonymous referee for their comments that improved the clarity of this work. This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI Australia) and the Pawsey Supercomputing Centre, NCRIS-enabled capabilities supported by the Australian Government, through award jh2. AA and MRK acknowledge support from the Australian Research Council through Laureate Fellowship FL220100020. MI wishes to thank HW Rix for lunchtime discussions in 2019 that motivated some of this work, and the Australia-Germany Joint Research Co-operation Scheme of Universities Australia for funding this travel.
@@
The codes and a subset of selected intermediate data used in this study are available at https://github.com/aruuniima/stellar-actions-I and https://www.mso.anu.edu.au/˜arunima/stellar-actions-I-data/. The initial simulation outputs and potential fits are not included in the public repository due to their size, but will be provided by the authors upon reasonable request.
Appendix A Comparison of epicyclic actions with more accurate action estimates
To assess the accuracy of the epicyclic approximation used throughout this work, we select a random sample of 10,000 stars from the final snapshot of our simulations and compare the actions we compute for these stars using the epicyclic approximation to those we obtain using a much more accurate (but considerably more computationally-intensive) action estimate based on the Stäckel approximation as implemented in the galpy package (Bovy, 2015). This method has been widely adopted in both observational and simulation-based studies of Galactic dynamics (e.g., Frankel et al., 2018; Trick et al., 2019; Jia et al., 2023; Palicio et al., 2025).
Our first step in this comparison is to construct the axisymmetric potential. Most previous studies have relied on the time-independent analytic Milky Way-like potentials provided by galpy. However, we have seen that our simulation, with its live stellar and gas disk, does not have an entirely constant potential. To remain consistent with the simulations, therefore, we instead extract the gravitational potential and acceleration fields directly from our simulation snapshot and input them into galpy as a custom potential. We do so as follows: first, we extract the positions, masses, and smoothing lengths of all particles – stars, gas, and dark matter – and use pytreegrav (Grudić & Gurvich, 2021) to compute the resulting gravitational potential and acceleration on a regular cylindrical grid in $(R,z,\phi)$ with dimensions $(40000,6000,16)$, covering the range $R\in[0.1\ \text{pc},40\ \text{kpc}]$, $z\in[-3,3]\ \text{kpc}$, and $\phi\in[0,2\pi]$. To ensure axisymmetry, we average the potential and both the radial and vertical accelerations over $\phi$, resulting in a 2D $(R,z)$ grid. As in our epicyclic approximation pipeline, we next smooth these grids to facilitate accurate numerical differentiation using the same smoothing approach as described in Appendix B for the potential used in the epicyclic approximation – see accompanying code linked in Data Availability for the full method. We then construct 2D spline fits for the potential, radial force, vertical force, and the $R$- and $z$-derivatives of the force, which are required for the numerical intgegrated used in galpy’s actionAngleStaeckel routine.
For each of our 10,000 sample stars, following galpy documentation, we integrate each star’s orbit in this potential for 50 internal time units (corresponding to $\approx 1.9$ Gyr) and use the resulting orbit data to estimate the local focal parameter $\Delta$ using the estimateDeltaStaeckel function. With the potential and $\Delta$ in hand, we then calculate actions using actionAngleStaeckel. This method successfully returns actions for $\approx 70\%$ of the sample. [^3] We have verified that the success rate is consistently around this value by repeating the process multiple times with different sample stars. The failures are due to galpy’s inability to estimate $\Delta$, likely stemming from numerical noise in the derivatives. However, since the action computation itself is generally robust to such errors, for these failed cases we simply fix $\Delta=0.5$, a typical value for disc orbits. With this substitution, the success rate increases to 98%.
Figure 16: The probability distribution of the squared relative error of epicyclic actions with respect to the more accurately determined actions for the 75% success rate sample (upper row) and the 98% success rate sample (lower row). The radial and vertical action errors are shown in blue and pink colours, respectively.
We compare the epicyclic and Stäckel-based action estimates in Figure 16, showing both the $\approx 70\%$ of the sample for which galpy works “out of the box” and the expanded sample using fixed $\Delta$. We find that the median squared relative errors in $J_{R}$ and $J_{z}$ for the 70% sample are 0.08 and 0.01, respectively, and 0.05 and 0.01, respectively, for the expanded sample. Errors in $J_{\phi}$ are negligible ($4\times 10^{-14}$ kpc km/s for all stars) and are not shown. These figures are consistent with the small differences we expect based on the findings of Solway et al. (2012), and are much smaller than the variations in mean-square action that we find in the main text. This confirms that the epicyclic approximation is sufficient for our purposes.
Appendix B Computing epicyclic frequencies
@@
In this appendix, we describe in detail the procedure that we use for calculating epicyclic frequencies $\kappa$ and $\nu$ for each star particle. We begin with the 2D potential grid that spans from 0.1 pc to 20 kpc in the radial direction and from 1-1 to +1+1 kpc in the vertical direction, computed as described in Section 3.1, and we must evaluate partial derivatives of this potential to compute $\kappa$ and $\nu$. Naive numerical evaluation of these derivatives yields extremely noisy estimates, and we must therefore adopt a smoothing procedure.
@@
The guiding radius $R_{g}$ and radial epicyclic frequency depend on $\partial\Phi/\partial R$ and $\partial^{2}\Phi/\partial R^{2}$ evaluated at $z=0$. Our first step to estimate these quantities is to extract the potential grid values within $\pm 2$ pc of the midplane, corresponding to five grid points in the vertical direction, and average them to produce a one-dimensional function of $R$. This averaging reduces the noise. We next apply a Gaussian filter [^4] This and all subsequent operations use the implementations provided in SciPy version 1.11.4; see Virtanen et al. (2020). with a standard deviation of 30 pc using ‘nearest’ mode to prevent artificial edge effects; this further smooths the data. Finally, we fit the smoothed data with a B-spline with smoothing parameter $s=200$ to obtain a continuous, differentiable function $\Phi(R,z=0)$.
@@
With this smooth function at hand, we are now prepared to evaluate the radial derivatives of $\Phi$ and the quantities that depend upon them. We first define a new radial grid with a resolution of 0.5 pc with its inner edge placed at the larger of 2 kpc and 5 kpc inward from the smallest radial position of any star at the time we are sampling, and its outer edge at the smaller of 17.5 kpc and 5 kpc outward from the largest stellar radial position; this ensures that we have good coverage over the full disc, but that we are not extrapolating to radii where we sample the potential poorly. We then evaluate $\Omega_{i}$ at each radial grid point $R_{i}$ from equation 5 using our B-spline representation of $\Phi(R)$. The corresponding specific angular momentum at each grid point is $L_{z,i}=R_{i}^{2}\Omega_{i}$, and thus our grid represents a set of $(R_{i},L_{z,i})$ pairs. Since the guiding radius $R_{g}$ for each star is defined implicitly by its angular momentum $L_{z,*}$ and the condition $L_{z,*}=R_{g}^{2}\Omega(R_{g})$, we can use the $(R_{i},L_{z,i})$ pairs as a lookup table to find the guiding radius corresponding to any $R_{g}$. In practice we do this by defining a new B-spline function $R_{g}(L_{z})$ from our $(R_{i},L_{z,i})$ pairs using a smoothing factor $s=2$, and use the resulting function to evaluate $R_{g}$ for every star. We then drop from our sample stars for which $R_{g}$ lies outside the range 2162-16 kpc, on the grounds that they lie too close to our grid edge to be reliable.
@@
The calculation of the vertical epicyclic frequency $\nu$ follows a similar but slightly more complex process, since we require the second derivative of $\Phi(z)$ at each star’s guiding radius (see equation 7), rather than for the radial case which only requires derivatives of $\Phi(R)$ evaluated at $z=0$. To achieve this, we construct a sparse radial grid with a resolution of 200 pc, spanning the range 2162-16 kpc to match the guiding radius constraints that we imposed earlier. At each radial grid point, we extract the potential from the 2D grid by averaging over all radial values within $\pm 1$ pc of the sparse grid point. This averaging, as in the radial case, reduces local noise and provides a better estimate of the galactic potential as a function of $z$. We then smooth the resulting $\Phi(z)$ data by applying a Savitzky-Golay filter with a window length of 101 pc and a polynomial order of 2, which smooths the data while preserving the underlying curvature, followed by a Gaussian filter with standard deviation of 2 pc, which we apply three times. Finally, we use the smoothed data as input to a B-spline fit with $s=50$ to generate a smooth, continuously-differentiable functional form of $\Phi(z)$ at each sparse radial grid point. We use this smooth function to evaluate $\partial^{2}\Phi/\partial z^{2}$ at $z=0$, the quantity required to evaluate $\nu$ using equation 7. This yields a set of $(R_{i},\nu_{i})$ pairs on our sparse radial grid. We fit these data with a B-spline fit with $s=1000$ to obtain a continuous function $\nu(R)$. Finally, we determine $\nu$ for each star by inputting its guiding radius into this function.